OpenStudy (anonymous):

Write the equation {y=mx+b} 1)Given:(6,4)&(4,8) 2)Given:(2,3)&(1,2)

OpenStudy (amistre64):

m is found by subtracting the points and stacking y/x

OpenStudy (amistre64):

(6,4) -(4,8) ------ 2,-4; m=-4/2 = -2 we can use this in: y = mx -mPx + Py , where (Px,Py) is either of the given points

OpenStudy (amistre64):

y = -2x--2(6)+4 y = -2x+12+4 y = -2x+16 if we use the other point: y = -2x--2(4)+8 y = -2x+8+8 y = -2x+16 same results